OpenStudy (anonymous):

differentiate function y=x^2 + 4x +3/(square root of x)

OpenStudy (anonymous):

2x+4 -(3/2)x^(-3/2)

OpenStudy (anonymous):

nothing is being multiplied so each is the power rule. ax^n = anx^(n-1) is the general formula for each where a is what is in front of x, x is the variable, n is the power

OpenStudy (anonymous):

Davidjohn wait a sec, you have to use the quotient rule here

OpenStudy (anonymous):

for the last one 3/x^1/2 is 3x^-1/2 \[3(-1/2)x^ ((-1/2)-(2/2))\]

OpenStudy (anonymous):

Oh it depends, is it y=x^2 + 4x +3/(square root of x) or y=(x^2 + 4x +3)/(square root of x) ?

OpenStudy (anonymous):

no there is no variable times a variable in the numerator or denominator. right if it were that then you would

OpenStudy (anonymous):

but i helped this person before and most were power rule so im assuming that this is too

OpenStudy (anonymous):

both of them have a parenthesis Andras

OpenStudy (anonymous):

So you need d/dx(x^2 + 4x +3)/(square root of x) ?

OpenStudy (anonymous):

but the top one has parenthasis in the numerator which would make the whole top term divided by the bottom rather than just the +3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok than you need to know that (f(x)/g(x))'={f'(x)*g(x) -g'(x)f(x)}/g^2(x)

OpenStudy (anonymous):

I guess david is doing it for you

OpenStudy (anonymous):

andras, the x^1/2 would be positive right? in the quotient rule

OpenStudy (anonymous):

The rule is there here f(x)=x^2 + 4x +3 and g(x)=sqrt(x)

OpenStudy (anonymous):

okay then, just making sure :) thanks

OpenStudy (anonymous):

So f'(x)=2x+4 g'(x)=1/2 x^(-1/2)

OpenStudy (anonymous):

yeah and the g(x)^2 would just be x

OpenStudy (anonymous):

so, ((2x + 4)(x^1/2)-(x^2 + 4x + 3)(1/2x^-1/2))/ (x)

OpenStudy (anonymous):

{(2x+4)(sqrtx)-(x^2+4x+3)(1/2x^(-1/2))}/x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now I guess this can be rearranged into a simple form

OpenStudy (anonymous):

haha true, but didnt want to do ALL the work ;)

OpenStudy (anonymous):

ok :-) Lets leave it than?

OpenStudy (anonymous):

Dany are u happy?

OpenStudy (anonymous):

okay you all confused me, so whats the answer o.O

OpenStudy (anonymous):

{(2x+4)(sqrtx)-(x^2+4x+3)(1/2x^(-1/2))}/x

OpenStudy (anonymous):

I need pen and paper to rearrange this into a simple form and I cannot be asked to search for one now :-)

OpenStudy (anonymous):

it all together, when you have a variable function divided by another function the rule is (f(x)/g(x))'={f'(x)*g(x) -g'(x)f(x)}/g^2(x) here f(x)=x^2 + 4x +3 and g(x)=sqrt(x) So f'(x)=2x+4 g'(x)=1/2 x^(-1/2) if put all together from andras and i, mostly the former ((2x + 4)(x^1/2)-(x^2 + 4x + 3)(1/2x^-1/2))/ (x) which i will simplify in a minute

OpenStudy (anonymous):

oh so i use the other rule, i was thinking i had to do the power rule.

OpenStudy (anonymous):

you would have to only if the whole numerator was not in parenthesis because then it would be just constants times a variable

OpenStudy (anonymous):

well in the paper it has the polynomial all divided by the square root of x

OpenStudy (anonymous):

y=x^2 + 4x +3/(square root of x) is power rule y=(x^2 + 4x +3)/(square root of x) is quotient rule

OpenStudy (anonymous):

but i dont think i have that rule yet in my notes in which we are doing this, all i have is the constant, power, and d/dx[cf(x)]=cf'(x)

OpenStudy (valpey):

\[y'(x) = -\frac{3}{2 x^{3/2}}+2 x+4\]

OpenStudy (anonymous):

that is what i put first but i guess the whole function is divided by the x^.5

OpenStudy (valpey):

\[y(x) = \frac{x^2 + 4x +3}{\sqrt{x}}\] \[y'(x) = \frac{3 x^2+4 x-3}{2 x^{3/2}}\]